OpenStudy (anonymous):

help with backwards intergral derrivate elvalate the following intergate sign y=( x^2+3x+5) dx

OpenStudy (anonymous):

use the rule dx=x+c the d/dx is 1x+3) DX

OpenStudy (anonymous):

so there fore we use x^2 dx = 1/n+1 x^n+1 +C right?

OpenStudy (anonymous):

yes right

OpenStudy (unklerhaukus):

is this your question?\[y=\int (x^2+3x+5)\text dx\]

OpenStudy (anonymous):

yes sir

OpenStudy (unklerhaukus):

yes apply \[\int ax^n=\frac{ax^{n+1}}{n+1}+c\]

OpenStudy (anonymous):

so do I use D/DX to apply to this or whole problem again that is where I am lost..

OpenStudy (unklerhaukus):

\[y=\int (x^2+3x+5)\text dx=\frac{x^{2+1}}{2+1}+\frac{3x^{1+1}}{1+1}+\frac {5x^{0+1}}{0+1}+{c}\]

OpenStudy (anonymous):

okay I thought you did d/DX and than use that

OpenStudy (anonymous):

not the problem without finding D/DX what am missing hrere??

OpenStudy (unklerhaukus):

i dont know what you mean

OpenStudy (anonymous):

I did the der first than tried to apply that form to it

OpenStudy (anonymous):

y=1x+3 D/DX

OpenStudy (unklerhaukus):

huh/? if \[f(x)=x^2+3x+5\]\[f'(x)=2x+3\], but i dont know why you are differentiating

OpenStudy (anonymous):

opps 2x not 1x. That is what I was asking do you need to differentate before using that formula

OpenStudy (unklerhaukus):

differentiating is kinda like the opposite to integrating, like + is oppositse to - and x is opposite to \(\div\),

OpenStudy (unklerhaukus):

if the question is an integration, differentiating wont help much

OpenStudy (unklerhaukus):

but once you have integrated , you could check if differentiating get you back to the original question,

OpenStudy (anonymous):

okay so the whole problem must used and be plugged in as such.

OpenStudy (anonymous):

for every x I use x^2 and so on.

OpenStudy (unklerhaukus):

for example \[y=\int10x\text dx=\frac{10x^2}{2}+c=5x^2+c\] \[\frac{\text dy}{\text dx}=\frac{\text d}{\text dx}(5x^2+c)=2\times5x=10x\] \[\text dy=10x\text dx\] \[\int\text dy=\int10x\text dx\]\[y=\int10x\text dx\]

OpenStudy (anonymous):

so pretty much you sub in all values for letters with in the formula and do the intergation from that point right?

OpenStudy (unklerhaukus):

what do you get?

OpenStudy (anonymous):

how do you use the equation function on this board??

OpenStudy (unklerhaukus):

i dont use the equation tool ive been typing the \(\LaTeX\) by hand for example type y=\int10x\text dx=\frac{10x^2}{2}+c=5x^2+c with \[\[ \]\] around the whole expression produces \[y=\int10x\text dx=\frac{10x^2}{2}+c=5x^2+c\]

OpenStudy (anonymous):

so in plugging in the rule i have x^2 dx=1/5+1 x^2 +c did sub right here?

OpenStudy (unklerhaukus):

i dont understand what you have done

OpenStudy (anonymous):

i sub in the values for what the formula called for

OpenStudy (anonymous):

what sucks here I have a TI 92 but no batteries until Wed or Th so I can't do this

OpenStudy (unklerhaukus):

\[y=\int (x^2+3x+5)\text dx\]\[=\int x^2\text dx+\int3x\text dx+\int5\]\[=\frac{x^{2+1}}{2+1}+\frac{3x^{1+1}}{1+1}+\frac {5x^{0+1}}{0+1}+{c}\]

OpenStudy (anonymous):

okay how do what goes where that is what I am missing here..

OpenStudy (unklerhaukus):

* there should be a dx after the 5

OpenStudy (anonymous):

there is.

OpenStudy (unklerhaukus):

you confuse me @godorovg

OpenStudy (anonymous):

I am asking how to know what is sub in and where when using the perfect power rule for intergrals

OpenStudy (unklerhaukus):

what substitution are you talking about ?

OpenStudy (anonymous):

you have the this rule x^2 dx=1/n+1 x^n+1 +c now I do I use this with y=x^2 +3x+5 that is my question.

OpenStudy (anonymous):

help with where 3x and the 5 goes with this and why?